OpenStudy (anonymous):

CaCO3(s) + H3PO4 (aq) -> Ca3(PO4)2(aq) + H20(l) + CO2(g) how many grams of phosphoric acid react with excess calcium carbonate to produce 3.74g ca3(po4)2

OpenStudy (aaronq):

After you have balanced the reaction, convert the mass given to moles. To interconvert between mass and moles, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles. -------------------------------------------------- Next, use the stoichiometric coefficients to find moles of reagent needed. Set up a ratio using the species of interest, like so: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) --------------------------------------------------------